The absolute visual magnitude of a star is 6.5 and for its temperature, the bolometric correction is − 0.3. Calculate the absolute bolometric magnitude and the luminosity of the star.
1
Expert's answer
2019-04-02T10:39:19-0400

The expression for the absolute magnitudes of a star are written as:


"M_{bol}=M_{v}+B.S. (1)"

where Mbol is the absolute bolometric magnitude of a star, Mv is the absolute visual magnitude of a star, B.C. is bolometric correction

Using (1) we got: Mbol = 6.5-0.3=6.2

The formula relating absolute bolometric magnitude with luminosity is as follows:


"L_{star}= L_ʘ\times 10^{-0.4M_{bol}} (2)"

where Lʘ is the zero-point luminosity (the luminosity of a star with Mbol=0)

We find:

Lstar=0.003 Lʘ


Answer:

Mbol = 6.2

Lstar=0.003 Lʘ